### v1

# Variational Gaussian Approximation in Replica Analysis of Parametric Models

Abstract
We revisit the replica method for analyzing inference and learning in parametric models, considering situations where the data-generating distribution is unknown or analytically intractable. Instead of assuming idealized distributions to carry out quenched averages analytically, we use a variational Gaussian approximation for the replicated system in grand canonical formalism in which the data average can be deferred and replaced by empirical averages, leading to stationarity conditions that adaptively determine the parameters of the trial Hamiltonian for each dataset. This approach clarifies how fluctuations affect information extraction and connects directly with the results of mathematical statistics or learning theory such as information criteria. As a concrete application, we analyze linear regression and derive learning curves. This includes cases with real-world datasets, where exact replica calculations are not feasible.
I Introduction
The statistical mechanics analysis of learning and inference from examples has a long history, starting with the seminal work of Seung, Somplinsky and Tishby in the early 1990s Sompolinsky et al. (1990); Seung et al. (1992). Since then, such approaches have become a standard topic in the statistical mechanics of disordered systems Engel (2001); Nishimori (2001). One of the central analytical tools in this line of research is the replica method Mezard et al. (1986); Charbonneau et al. (2023); Montanari and Sen (2024). It provides a powerful method to evaluate the average behavior of statistical quantities such as free energy and prediction error. Unlike classical learning theory, which often yields only bounds for those quntities Shalev-Shwartz and Ben-David (2014), the replica method can give sharp predictions.
Despite its strength, the replica method has a major limitation. To take quenched averages analytically, one usually has to assume highly idealized data-generating distributions. This simplification is required even in problems as basic as the standard linear regression. Hence, the replica method has been regarded as a mathematical technique for guessing exact solutions under highly restricted conditions, rather than a tool for predicting universal structures of learning and inference, or for describing the behavior of them in realistic data, although some exceptions exist Malzahn and Opper (2000, 2001, 2002, 2003, 2005); Yoshino (2018); Loureiro et al. (2021); Yoshino (2020, 2023).
A particularly notable line of work in the context of learning and inference was provided by the series of studies by Malzahn and Opper Malzahn and Opper (2000, 2001, 2002, 2003, 2005). By combining the replica method with variational approximations, they successfully obtained an approximate solution to learning curves for Gaussian process regression and hard-margin support vector machines without assuming a concrete data distributions. The use of the variational approximation simultaneously offered two important advantages. They relax the dependence on overly simplified data distributions and, moreover, it allowed a systematic way of giving approximate predictions even when exact replica calculations were not trivial. These advantages may be useful in the analysis of learning with modern complex models.
However, their analysis was mainly restricted to non-parametric models, and did not address parameter-space formulations that do not necessarily reduce to Gaussian process analyses, which may be more directly relevant to parametric models such as modern neural networks. In this work, we investigate the use of variational Gaussian approximation in replica analysis of parametric models. As a first step, we represent general asymptotic properties as well as application to linear regression. While most of the resulting formulae are already known from mathematical statistics and learning theory, this study represents a methodological step toward relaxing the restricted assumptions of conventional replica analysis of parametric models and may broaden their applicability to a wider class of learning and inference problems.
The remainder of the paper is organized as follows. In Section II, we describe the setup of our analysis, focusing on independent and identically distributed (i.i.d.) data and parameter inference based on the Boltzmann distribution, also known as the Gibbs posterior Catoni (2007); Bissiri et al. (2016). Section III introduces the framework of replica method for learning and inference. In Section IV, we formulate the variational Gaussian approximation. Section V discusses general asymptotic properties. In Section VI, we apply the method to linear regression as a concrete example. Finally, Section VII summarizes the results and outlines possible directions for future research.
II Setup
We consider a dataset $D_{n}=\{z_{i}\}_{i=1}^{n}$ of $n$ i.i.d. samples drawn from an unknown distribution $p_{\rm{data}}$ defined on a sample space $\mathcal{Z}$}. Each data point $z_{i}$ represents a generic observation. In unsupervised learning it can be a feature vector $x_{i}\in\mathbb{R}^{d}$, and in supervised learning it corresponds to an input–output pair $(x_{i},y_{i})$.
In order to describe learning and inference with a parametric model, we introduce a Boltzmann distribution, also known as the Gibbs posterior Catoni (2007); Bissiri et al. (2016), on the parameter space $\Theta$:
$p^{\beta}(\theta\mid D_{n})=\frac{1}{Z^{\beta}(D_{n})}e^{-\beta\,{\mathcal{L}}(\theta;D_{n})},$ (1)
where
${\mathcal{L}}(\theta;D_{n})=\sum_{i=1}^{n}l(\theta;z_{i})+\frac{\lambda}{2}\|\theta\|_{2}^{2},$ (2)
and the parameter $\theta\in\Theta$ is represented as an $N$-dimensional vector $\theta=(\theta_{1},\ldots,\theta_{N})$.
The function $l(\theta;z_{i})$ represents a loss associated with each data point, such as a negative log-likelihood, and the term $\tfrac{\lambda}{2}\|\theta\|_{2}^{2}$ corresponds to the weight decay. The sum $\mathcal{L}$} corresponds to a Hamiltonian in statistical mechanics. More generally, regularization can be incorporated into the measure over $\theta$, which would be more suitable when the regularization is more complicated, or the parameter is discrete. The parameter $\beta>0$ is an inverse temperature. The case $\beta=1$ corresponds to Bayesian inference with the posterior distribution, while in the limit $\beta\to\infty$ the distribution concentrates on the minimizers of ${\mathcal{L}}(\theta;D_{n})$, which corresponds to empirical risk minimization.
From this distribution we can consider a random variable $\hat{\theta}(D_{n})$ sampled as
$\hat{\theta}(D_{n})\sim p^{\beta}(\cdot\mid D_{n}).$ (3)
The aim of statistical mechanics analysis is to study how $\hat{\theta}(D_{n})$ fluctuates according to thermal noise and quenched randomness, and how the statistical averages with respect to this distribution behave. In this way one can predict the behavior of quantities of interest in inference and learning, such as the prediction error.
We are particularly interested in the training error $\epsilon_{\rm tr}$, the prediction error $\epsilon_{\rm pred}$, and their difference, the generalization gap $\delta\epsilon$, which are defined through an error function $\epsilon(\cdot;\cdot):\Theta\times{\mathcal{Z}}\to\mathbb{R}$ as
$\epsilon_{\rm tr}(D_{n})$ $=\frac{1}{n}\sum_{i=1}^{n}\langle\epsilon(\theta;z_{i})\rangle^{\beta,n}$ (4)
$\epsilon_{\rm pred}(D_{n})$ $=\mathbb{E}_{z\sim p_{\rm{data}}}\left[\langle\epsilon(\theta;z)\rangle^{\beta,n}\right]$ (5)
$\delta\epsilon(D_{n})$ $=\epsilon_{\rm pred}(D_{n})-\epsilon_{\rm tr}(D_{n}).$ (6)
Here $\langle\;\cdot\;\rangle^{\beta,n}$ denotes the average with respect to the Boltzmann distribution  (1). These quantities depend on the training dataset $D_{n}$ and are therefore random. It is often of interest to study their average behavior, such as
$\bar{\epsilon}_{\rm tr}$ $=\mathbb{E}_{D_{n}}\left[\epsilon_{\rm tr}(D_{n})\right]$ (7)
$\bar{\epsilon}_{\rm pred}$ $=\mathbb{E}_{D_{n}}\left[\epsilon_{\rm pred}(D_{n})\right]$ (8)
$\bar{\delta\epsilon}$ $=\mathbb{E}_{D_{n}}\left[\delta\epsilon(D_{n})\right].$ (9)
We also remark that the error metric $\epsilon$ does not have to coincide with the loss function $l$ used in the definition of $\mathcal{L}$}.
III Replica Method
To systematically investigate the fluctuations of $\hat{\theta}(D_{n})$ defined in  (3) with respect to thermal noise and quenched randomness, it is useful to consider the replicated system defined for natural numbers $r=1,2,\ldots$ as a density on $\Theta^{r}$:
$p_{r}^{\beta,n}(\bm{\theta})=\frac{1}{\Xi_{r}^{\beta,n}}\,\mathbb{E}_{D_{n}}\left[\prod_{a=1}^{r}e^{-\beta{\mathcal{L}}(\theta^{a};D_{n})}\right],\quad\theta^{a}\in\Theta,a\in[r],$ (10)
where $\Xi_{r}^{\beta,n}$ is normalization constant and $r]\equiv\{1,2,\ldots,r\}$. We also use the shorthand notation $\bm{\theta}\in\Theta^{r}$ to denote the concatenated vector of replicas, that is, the long vector obtained by stacking $\theta^{1},\ldots,\theta^{r}$ vertically:
$\bm{\theta}=\begin{pmatrix}\theta^{1}\\ \theta^{2}\\ \vdots\\ \theta^{r}\end{pmatrix}$. (11)
Since the definition already involves an average over the dataset $D_{n}$, this distribution is no longer conditioned on $D_{n}$. However, the fluctuation of $\hat{\theta}(D_{n})$ with respect to quenched randomness is encoded in the correlations among replicas, which can be accessed through the correlation functions such as $\mathbb{E}_{r}^{\beta,n}[\theta_{i}^{1}\theta_{i}^{2}]$, where $\mathbb{E}_{r}^{\beta,n}[\cdot]$ denotes the average with respect to the replicated system  (10). By extrapolating these quantities to real values of $r$, one can obtain the desired information.
III.1 Grand canonical formalism
For subsequent analysis, it is convenient to rewrite the replicated system  (10) as follows. The expectation with respect to $D_{n}$ is taken over i.i.d. sampling $z_{i}\sim p_{\rm{data}},\,i\in[n]$. In the i.i.d. case, this averaging is equivalent to sampling $n$ data points with replacement from a sufficiently large dataset $\tilde{D}=\{\tilde{z}_{i}\}_{i=1}^{\tilde{n}},\,\tilde{z}_{i}\sim p_{\rm{data}}$, since the empirical distribution of a sufficiently large dataset converges to the true distribution $p_{\rm{data}}$ and can approximate it with arbitrary accuracy. This is in the same spirit as the bootstrap method developed by Efron Efron (1979); Efron and Tibshirani (1994). Let $\tilde{c}_{i}\in[\tilde{n}]\cup\{0\}$ denote the number of times $\tilde{z}_{i}$ is sampled. Then
$\mathbb{E}_{D_{n}}\left[\prod_{a=1}^{r}e^{-\beta{\mathcal{L}}(\theta^{a};D_{n})}\right]\simeq\mathbb{E}_{\tilde{\bm{c}}}\left[\prod_{i=1}^{\tilde{n}}e^{-\tilde{c}_{i}\beta\sum_{a=1}^{r}l(\theta^{a};\tilde{z}_{i})}\right]e^{-\tfrac{\beta\lambda}{2}\sum_{a=1}^{r}\|\theta^{a}\|_{2}^{2}},$ (12)
[…]
$p_{r,{\rm GC}}^{\beta,n}(\theta^{1},\ldots,\theta^{r})$ $=\frac{1}{\Xi_{r,{\rm GC}}^{\beta,n}}e^{nH(\bm{\theta})},$ (13)
$H(\bm{\theta})$ $=\mathbb{E}_{z\sim p_{\rm{data}}}\left[e^{-\beta\sum_{a=1}^{r}l(\theta^{a};z)}\right]-\tfrac{\beta\lambda}{2n}\sum_{a=1}^{r}\|\theta^{a}\|_{2}^{2}.$ (14)
From the viewpoint of statistical mechanics, this corresponds to the grand canonical ensemble where the number of data points, instead of the number of particles, fluctuates with chemical potential $\beta^{-1}\log n$. Hence, we refer to  (13) as the grand canonical replicated system. Appendix A presents a derivation more directly emphasizing the formal analogy with the grand canonical ensemble. When computing averages, either  (10) or  (13) can be used in principle, but as we shall see, the GC formalism often simplifies the analysis.
III.2 General formulae for errors
Before proceeding to the variational approximation of the replicated system, we here summarize general formulas for error metrics  (7)– (9) based on the replica method. For this, we rewrite the normalization factor of the Boltzmann distribution, which appears in the definitions of the error metrics, as $1/Z^{\beta,n}(D_{n})=\lim_{r\to 0}(Z^{\beta,n}(D_{n}))^{r-1}$, and note that the normalization constant of the replicated system converges to one as $r\to 0$. As is common in replica calculations, after symmetrization over replica indices the error metrics can be expressed as
$\bar{\epsilon}_{\rm tr}$ $=\lim_{r\to 0}\lim_{\gamma\to 0}\frac{d}{d\gamma}\log\mathbb{E}_{D_{n}}\left[\int e^{-\beta\sum_{a=1}^{r}{\mathcal{L}}(\theta^{a};D_{n})+\frac{\gamma}{nr}\sum_{i=1}^{n}\sum_{a=1}^{r}\epsilon(\theta^{a};z_{i})}d\theta^{1}\ldots d\theta^{r}\right],$ (15)
$\bar{\epsilon}_{\rm pred}$ $=\lim_{r\to 0}\mathbb{E}_{z\sim p_{\rm{data}}}\left[\mathbb{E}_{r}^{\beta,n}\left[\frac{1}{r}\sum_{a=1}^{r}\epsilon(\theta^{a};z)\right]\right],$ (16)
where $\lim_{r\to 0}$ should be interpreted in the sense of the replica trick, as the extrapolation of results from integer $r$ to zero. Recall that $\mathbb{E}_{r}^{\beta,n}[\cdot]$ is the expectation with respect to the replicated system  (10).
By rewriting the dataset average as in the derivation of the GC formalism, and replacing the expectation $\mathbb{E}_{r}^{\beta,n}$ with $\mathbb{E}_{r,{\rm GC}}^{\beta,n}$, we obtain the following compact forms:
$\bar{\epsilon}_{\rm tr}$ $=\mathbb{E}_{z\sim p_{\rm{data}}}\left[\lim_{r\to 0}\frac{1}{r}\sum_{a=1}^{r}\mathbb{E}_{r,{\rm GC}}^{\beta,n}\left[\epsilon(\theta^{a};z)e^{-\beta\sum_{b=1}^{r}l(\theta^{b};z)}\right]\right],$ (17)
$\bar{\epsilon}_{\rm pred}$ $=\mathbb{E}_{z\sim p_{\rm{data}}}\left[\lim_{r\to 0}\frac{1}{r}\sum_{a=1}^{r}\mathbb{E}_{r,{\rm GC}}^{\beta,n}\left[\epsilon(\theta^{a};z)\right]\right],$ (18)
which further yield
$\bar{\delta\epsilon}=\mathbb{E}_{z\sim p_{\rm{data}}}\left[\lim_{r\to 0}\frac{1}{r}\sum_{a=1}^{r}\mathbb{E}_{r,{\rm GC}}^{\beta,n}\left[(1-e^{-\beta\sum_{b=1}^{r}l(\theta^{b};z)})\epsilon(\theta^{a};z)\right]\right].$ (19)
In these expressions, the expectation with respect to the GC replicated system plays the role of a cavity bias at the data point $z$. In particular, the formula for the training error shows that the factor $e^{-\beta\sum_{a=1}^{r}l(\theta^{a};z)}$ appears as a bias because the evaluation point is included in the training dataset.
III.2.1 Remark (link to PCIC/WAIC)
Here we briefly note a connection with information criteria, which are estimators for the generalization gap Konishi and Kitagawa (2008). If we expand the factor $e^{-\beta\sum_{a=1}^{r}l(\theta^{a};z)}$ in a Taylor series and analytically continue as $r\to 0$ without considering replica symmetry breaking fields, we obtain
$\bar{\delta\epsilon}=\mathbb{E}_{z\sim p_{\rm{data}}}\Bigl{[}\,\beta\left(\langle l(\theta;z)\epsilon(\theta;z)\rangle^{\beta,n}-\langle l(\theta;z)\rangle^{\beta,n}\langle\epsilon(\theta;z)\rangle^{\beta,n}\right)\Bigr{]}+\ldots.$ (20)
If the expectation over $z$ is replaced by the empirical average over the training data, the first-order term coincides with the posterior covariance information criterion (PCIC) Iba and Yano (2023), and for $l=\epsilon$ it coincides with widely applicable information criterion (WAIC) Watanabe (2013). In particular, when $l=\epsilon$, the expansion corresponds to a cumulant expansion with respect to $\beta l$.
Unfortunately, it is difficult to obtain useful bounds on the higher-order terms from this representation, because simple bounds such as $1-e^{-x}\leq x$, valid for $r\in\mathbb{N}$, are not guaranteed to yield meaningful bounds after analytic continuation to $r\to 0$. Therefore, while this correspondence is interesting, its practical usefulness remains unclear to the author at present.
IV Variational Gaussian Approximation (VGA)
The replicated systems  (10) and  (13) are in general intractable unless the data generating distribution $p_{\rm{data}}$ is simple enough. To proceed, we adopt a variational approximation in GC formalism. We introduce a trial Hamiltonian $\tilde{H}$, which is determined by the stationarity conditions of the variational free energy
${\mathcal{F}}_{r,{\rm GC}}^{\beta,n}=\log\tilde{\Xi}_{r,{\rm GC}}^{\beta,n}+n\,\tilde{\mathbb{E}}_{r,{\rm GC}}^{\beta,n}\left[H(\bm{\theta})-\tilde{H}(\bm{\theta})\right],$ (21)
where $\tilde{\mathbb{E}}_{r,{\rm GC}}^{\beta,n}[\cdot]$ denotes the average with respect to the approximate replicated system proportional to $e^{n\tilde{H}(\bm{\theta})}$. The derivation of this variational free energy follows the standard perturbative variational principle Parisi (1988) and, for integer $r$, it provides a bound on the true free energy. However, in the limit $r\to 0$, its leading term $\lim_{r\to 0}r^{-1}{\mathcal{F}}_{r,{\rm GC}}^{\beta,n}$ is not necessarily guaranteed to bound the true free energy $\mathbb{E}_{D_{n}}[\log Z^{\beta,n}(D_{n})]$. Hence, instead of naively maximizing or minimizing with respect to the parameters of the trial Hamiltonian, we have to consider stationary conditions. Appendix B illustrates a simple example in which the variational parameters are at a saddle point rather than a maximum or minimum. It should be noted that the stationarity conditions are considered solely from the requirement of choosing the trial Hamiltonian optimally. Hence, unlike in many derivations of exact solutions, taking the thermodynamic limit is not necessary here, although one could consider an appropriate limit later.
The simplest trial Hamiltonian that incorporates interactions between parameters is a quadratic form:
$\tilde{H}(\bm{\theta})=-\frac{1}{2}(\bm{\theta}-\bm{m})^{\top}Q^{-1}(\bm{\theta}-\bm{m}),$ (22)
where $\bm{m}=[m_{i}^{a}]_{i\in[N],\,a\in[r]}\in\mathbb{R}^{Nr}$ is the concatenated vector formed by stacking $m^{a}$, and $Q=[Q_{ij}^{ab}]_{i,j\,\in\,[N],\,a,b\,\in\,[r]}\in\mathbb{R}^{Nr\times Nr}$ is a covariance matrix. We remark that, unlike in the conventional spin-glass literature where $Q$ typically denotes second moments (the so-called overlaps), here $Q$ corresponds to the covariance matrix. For now, we consider a general structure of the mean and covariance, though more specific assumptions may be adopted depending on prior knowledge or computational constraints.
With the trial Hamiltonian  (22), the variational free energy takes the following form, up to an additive constant,
${\mathcal{F}}_{r,{\rm GC}}^{\beta,n}$ $={\mathcal{F}}_{\rm ent}+{\mathcal{F}}_{\rm reg}+{\mathcal{F}}_{\rm int},$ (23)
${\mathcal{F}}_{\rm ent}$ $=\frac{1}{2}\log\det(Q/n),$ (24)
${\mathcal{F}}_{\rm reg}$ $=-\frac{\beta\lambda}{2n}\sum_{a=1}^{r}\sum_{i=1}^{N}\left((m_{i}^{a})^{2}+Q_{ii}^{aa}\right),$ (25)
${\mathcal{F}}_{\rm int}$ $=n\,\mathbb{E}_{z\sim p_{\rm{data}}}\left[\tilde{\mathbb{E}}_{r,{\rm GC}}^{\beta,n}\left[e^{-\beta\sum_{a=1}^{r}l(\theta^{a};z)}\right]\right].$ (26)
In the following, the variational parameters $\bm{m}$ and $Q$ are determined under the replica-symmetric (RS) assumption by imposing the stationarity conditions of  (23). We then use the approximate distribution $e^{n\tilde{H}(\bm{\theta})}/\tilde{\Xi}_{r,{\rm GC}}^{\beta,n}$ instead of the true replicated system  (13) to compute relevant quantities. For example, the generalization gap can be evaluated as
$\bar{\delta\epsilon}\simeq\mathbb{E}_{z\sim p_{\rm{data}}}\left[\lim_{r\to 0}\tilde{\mathbb{E}}_{r,{\rm GC}}^{\beta,n}\left[\frac{1}{r}\sum_{a=1}^{r}\epsilon(\theta^{a};z)\,\bigl{(}1-e^{-\beta\sum_{b=1}^{r}l(\theta^{b};z)}\bigr{)}\right]\right].$ (27)
It is worth emphasizing that the use of the grand canonical formulation yields the interaction term ${\mathcal{F}}_{\rm int}$ in a form where the expectation over the data distribution $z\sim p_{\rm{data}}$ is taken after averaging the factor $e^{-\beta\sum_{a=1}^{r}l(\theta^{a};z)}$ with respect to the approximate distribution. If we instead use the original formulation  (10), one would have to deal with the term $\log\mathbb{E}_{z\sim p_{\rm{data}}}[e^{-\beta\sum_{a=1}^{r}l(\theta^{a};z)}]$ inside the average with respect to the approximate distribution. Such a factor is difficult to handle unless the expectation over $p_{\rm{data}}$ can be computed analytically. This makes progress almost impossible unless the average over $p_{\rm{data}}$ is analytically tractable. By contrast, in the present formulation one may obtain a tractable expression for the interaction term by averaging with respect to the approximate distribution, provided that the trial Hamiltonian is sufficiently simple.
In practice, we derive the stationarity conditions while keeping the expectation over $p_{\rm{data}}$ explicit, and when needed we replace it by the empirical average. This makes it possible to obtain approximate formulae that apply to general datasets.
V General Results
In this section, we summarize general consequences of the RS parameterization that are largely independent of the specific form of the loss function $l$.
We consider the RS parameterization
$Q_{ij}^{ab}$ $=q_{ij}+\delta_{ab}\frac{\chi_{ij}}{\beta},$ (28)
$m_{i}^{a}$ $=m_{i}.$ (29)
This is equivalent to assuming that the estimator $\hat{\theta}(D_{n})$ behaves as an effective random variable of the form
$\begin{split}&\hat{\theta}(D_{n})=m+\xi+\eta,\\ &\xi\sim{\mathcal{N}}(0,q/n),\quad\eta\sim{\mathcal{N}}(0,\chi/(\beta n)),\end{split}$ (30)
where $\xi$ represents quenched fluctuations and $\eta$ represents thermal fluctuations. Also, $m$ is a parameter that explains how the training data breaks the symmetry of the model parameter.
Conditioned on $\xi\sim{\mathcal{N}}(0,q/n)$ and $z\sim p_{\rm{data}}$, we introduce the auxiliary densities on $\Theta$:
$\phi(\theta\mid\xi,z)$ $=\frac{1}{Z_{\phi}}e^{-\frac{\beta n}{2}(\theta-m-\xi)^{\top}\chi^{-1}(\theta-m-\xi)-\beta l(\theta;z)},$ (31)
$\phi_{\backslash z}(\theta\mid\xi)$ $=\frac{1}{Z_{\phi_{\backslash z}}}e^{-\frac{\beta n}{2}(\theta-m-\xi)^{\top}\chi^{-1}(\theta-m-\xi)}.$ (32)
[…]
For any function $g:\Theta\times{\mathcal{Z}}\to\mathbb{R}$, apart from the difference between the factors n1n-1 and $n$, we obtain
$\mathbb{E}_{D_{n}}\left[g(\langle\theta\rangle^{\beta,n},z_{i})\right]\simeq\mathbb{E}_{z\sim p_{\rm{data}},\,\xi\sim{\mathcal{N}}(0,q/n)}\left[g(\langle\theta\rangle_{\phi},z)\right],$ (33)
for $z_{i}\in D_{n}$, while for a fresh sample $\tilde{z}\sim p_{\rm{data}}$ independent of $D_{n}$,
[…]
In the zero-temperature limit $\beta\to\infty$, these reduce to
$\mathbb{E}_{D_{n}}\left[g(\langle\theta\rangle^{\beta,n},z_{i})\right]$ $\to\mathbb{E}_{z,\xi}\left[g(\theta^{\ast},z)\right],$ (35)
[…]
where
$\theta^{\ast}=\arg\min_{\theta}\frac{n}{2}(\theta-m-\xi)^{\top}\chi^{-1}(\theta-m-\xi)+l(\theta;z).$ (37)
Consequently, the averaged generalization gap takes the compact form
[…]
$\overset{\beta\to\infty}{\longrightarrow}\mathbb{E}_{z,\xi}\left[\epsilon(m+\xi;z)-\epsilon(\theta^{\ast};z)\right].$ (39)
V.1 Stationarity conditions
The variational parameters $m,q,\chi$ are determined by the following stationarity conditions:
$0$ $=\lambda m+n\mathbb{E}_{z\sim p_{\rm{data}}}\left[\langle\nabla_{\theta}l(\theta;z)\rangle_{\phi}\right],$ (40)
$\chi^{-1}q\chi^{-1}$ $=\mathbb{E}_{z\sim p_{\rm{data}}}\left[\langle\nabla_{\theta}l(\theta;z)\rangle_{\phi}\langle\nabla_{\theta}l(\theta;z)\rangle_{\phi}^{\top}\right],$ (41)
$\chi^{-1}$ $=\frac{\lambda}{n}I_{N}+\mathbb{E}_{z\sim p_{\rm{data}}}\left[\left.\nabla_{\gamma}\langle\nabla_{\theta}l(\theta+\gamma;z)\rangle_{\phi_{\gamma}}^{\top}\right|_{\gamma=0}\right],$ (42)
where
$\phi_{\gamma}(\theta\mid\xi,z)\propto e^{-\frac{\beta n}{2}(\theta-m-\xi)^{\top}\chi^{-1}(\theta-m-\xi)-\beta l(\theta+\gamma;z)}.$ (43)
These conditions correspond, respectively, to the stationarity of the population loss gradient, a (non-centered) covariance-like quantity for the gradients, and a regularized Hessian-like quantity. They are, however, evaluated under thermal averages, and therefore do not coincide with naive population quantities at finite $n$.
It may be useful to comment on the stationarity condition for $m$. Suppose that in the condition  (40) both $q$ and $\chi$ vanish. In this case, the stationary condition reduces to $0=\lambda m+n\,\mathbb{E}_{z}[\nabla_{m}l(m;z)]$, which corresponds to the stationary point of the population loss, i.e., an ideal estimator. However, in general, this is of course not the case. Hence, one important question is how this ideal condition is modified by the thermal fluctuations $\chi$ and the quenched randomness $q$. To reveal the structure of the solution, one typically needs either simplifying assumptions or numerical analysis, yet examining the form of the governing equations already provides insight into how well the estimator learns the average direction encoded by $m$. We will return to this point in subsubsection VI.2.1.
V.2 Asymptotic properties
We now examine the asymptotic properties at $n\to\infty$. To simplify the discussion, we consider $\beta\to\infty$ and assume that the Hessian of the population loss at the stationary point is positive definite. Considering the perturbative expansion of $\theta^{\ast}$ in $1/n$ as
$\theta^{\ast}=\theta_{0}^{\ast}+\frac{1}{n}\theta_{1}^{\ast}+\cdots,$ (44)
and straightforwardly expanding the stationary condition for the optimization problem  (37) $n\chi^{-1}(\theta^{\ast}-m-\xi)+\nabla_{\theta}l(\theta^{\ast};z)=0,$ we obtain
$\theta_{0}^{\ast}=0,\qquad\theta_{1}^{\ast}=\chi_{0}^{-1}\nabla_{\theta}l(m_{0};z),$ (45)
where $m_{0}$ satisfies the stationary condition of the population loss
$0=\mathbb{E}_{z\sim p_{\rm{data}}}\left[\nabla_{\theta}l(m_{0};z)\right],$ (46)
and $\chi_{0}$ denotes the inverse Hessian of the population loss at $m_{0}$:
$\chi_{0}=\left(\mathbb{E}_{z\sim p_{\rm{data}}}[\nabla_{\theta}l(m_{0};z)\nabla_{\theta}l(m_{0};z)^{\top}]\right)^{-1}.$ (47)
In this regime, the generalization gap is given by
$\bar{\delta\epsilon}=\frac{1}{n}\operatorname{Tr}\left[\chi_{0}\,\mathbb{E}_{z\sim p_{\rm{data}}}[\nabla_{\theta}\epsilon(m_{0};z)\nabla_{\theta}l(m_{0};z)^{\top}]\right]+{\mathcal{O}}(n^{-2}),$ (48)
which is precisely the Takeuchi Information Criterion Takeuchi (1976); Seung et al. (1992); Amari and Murata (1993); Amari et al. (1995). If the regularization parameter scales as $\lambda n$, then the condition for $m_{0}$ and $\chi_{0}^{-1}$ should be replaced by
$0$ $=\lambda m_{0}+\mathbb{E}_{z\sim p_{\rm{data}}}[\langle\nabla_{\theta}l(m_{0};z)\rangle],$ (49)
$\chi_{0}^{-1}$ $=\lambda I_{N}+\mathbb{E}_{z\sim p_{\rm{data}}}[\nabla_{\theta}l(m_{0};z)\nabla_{\theta}l(m_{0};z)^{\top}].$ (50)
VI Application to Linear Models
As a concrete example, we consider linear regression. In this setting, the data consist of input–output pairs $(x_{i},y_{i})$ with $x_{i}\in\mathbb{R}^{d}$ and $y_{i}\in\mathbb{R}$, and the parameter vector corresponds to the regression coefficients. For clarity, we denote the parameter by $w\in\mathbb{R}^{d}$ instead of the generic notation $\theta$. In linear regression, the input dimension coincides with the parameter dimension, so that $N=d$. For notational convenience, we explicitly indicate that the variational parameters are associated with $w$, writing them as $m_{w}$ and $Q_{w}$. With a slight abuse of notation, we replace the general loss $l(\theta;z)$ by a function $l:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ that depends on the response $y$ and the scaled inner product $\langle x,w\rangle\equiv d^{-1/2}\sum_{i=1}^{d}x_{i}w_{i},$ so that in the linear regression setting the loss takes the form $l(y,\langle x,w\rangle)$. To focus on the usual ridge-regularized empirical risk minimization, it is sufficient to consider the limit $\beta\to\infty$.
VI.1 General Properties
We now consider the general properties of linear models. Under the RS parametrization $m_{w}^{a}=m_{w},\quad Q_{w}^{ab}=q_{w}+\delta_{ab}\chi_{w}/\beta,$ the output of the model $\hat{f}(x)=\langle x,\hat{w}(D_{n})\rangle$ with $\hat{w}(D_{n})\sim p^{\beta}(\cdot\mid D_{n})$ can be effectively represented as the random variable
$\hat{f}(x)=m_{f}(x)+\xi_{f}+\eta_{f},$ (51)
where
$m_{f}(x)$ $=\langle x,m_{w}\rangle,$ (52)
$\xi_{f}$ $\sim{\mathcal{N}}(0,q_{f}(x)),$ (53)
$\eta_{f}$ $\sim{\mathcal{N}}(0,\chi_{f}(x)/\beta).$ (54)
Here, $\xi_{f}$ accounts for quenched randomness and $\eta_{f}$ for thermal fluctuations, with the variances given by
$q_{f}(x)$ $=\frac{1}{dn}\sum_{i,j=1}^{d}x_{i}x_{j}q_{w,ij},$ (55)
$\chi_{f}(x)$ $=\frac{1}{dn}\sum_{i,j=1}^{d}x_{i}x_{j}\chi_{w,ij}.$ (56)
Thus, the model output is characterized by a Gaussian process whose mean and variance both depend on the input $x$.
Similar to  (31)– (34), we can also describe the effective fluctuations of observables at training samples $z_{i}\in D_{n}$ and at fresh samples $\tilde{z}\sim p_{\rm{data}}$ not contained in $D_{n}$. Conditioned on $\xi_{f}$ and $z=(x,y)$, we introduce the distributions
$\phi^{\rm lin}(f\mid\xi_{f},z)$ $=\frac{1}{Z_{\phi}^{\rm lin}}e^{-\frac{\beta(f-m_{f}(x)-\xi_{f})^{2}}{2\chi_{f}(x)}-\beta l(y,f)},$ (57)
$\phi^{\rm lin}_{\backslash z}(f\mid\xi_{f},x)$ $=\frac{1}{Z_{\phi_{\backslash z}}^{\rm lin}}e^{-\frac{\beta(f-m_{f}(x)-\xi_{f})^{2}}{2\chi_{f}(x)}}.$ (58)
[…]
Using these distributions, for any function $g:\mathbb{R}\times{\mathcal{Z}}\to\mathbb{R}$, we then obtain
$\mathbb{E}_{D_{n}}\left[g\left(\bigl{\langle}\langle x_{i},w\rangle\bigr{\rangle}^{\beta,n},z_{i}\right)\right]$ $=\mathbb{E}_{z}\left[\mathbb{E}_{\xi_{f}\sim{\mathcal{N}}(0,q_{f}(x))}\left[g\bigl{(}\langle f\rangle_{\phi^{\rm lin}},z\bigr{)}\right]\right]$ (59)
$\overset{\beta\to\infty}{\longrightarrow}\mathbb{E}_{z}\left[\mathbb{E}_{\xi_{f}\sim{\mathcal{N}}(0,q_{f}(x))}\left[g(f^{\ast},z)\right]\right],$ (60)
[…]
$\overset{\beta\to\infty}{\longrightarrow}\mathbb{E}_{z}\left[\mathbb{E}_{\xi_{f}\sim{\mathcal{N}}(0,q_{f}(x))}\left[g\bigl{(}m_{f}(x)+\xi_{f},z\bigr{)}\right]\right],$ (62)
where
$f^{\ast}=\arg\min_{f}\frac{(f-m_{f}(x)-\xi_{f})^{2}}{2\chi_{f}(x)}+l(y,f).$ (63)
Accordingly, if the error is measured by a function $\epsilon(\cdot,\cdot):\mathbb{R}\times\mathbb{R}\to\mathbb{R}$, the data-averaged generalization gap takes the form
[…]
$\to\mathbb{E}_{z\sim p_{\rm{data}}}\left[\mathbb{E}_{\xi_{f}\sim{\mathcal{N}}(0,q_{f}(x))}\left[\epsilon(y,m_{f}(x)+\xi_{f})-\epsilon(y,f^{\ast})\right]\right].$ (65)
In the limit $\beta\to\infty$, the stationarity conditions take the following form, where $l^{\prime}$ denotes the derivative of $l$ with respect to its second argument:
$0$ $=\lambda m_{w}+n\,\mathbb{E}_{z,\xi_{f}}\left[\frac{x}{\sqrt{d}}\,l^{\prime}\bigl{(}y,f^{\ast}(x)\bigr{)}\right],$ (66)
$\chi_{w}^{-1}q_{w}\chi_{w}^{-1}$ $=\mathbb{E}_{z,\xi_{f}}\left[\frac{xx^{\top}}{d}\,\bigl{(}l^{\prime}(y,f^{\ast}(x))\bigr{)}^{2}\right],$ (67)
$\chi_{w}^{-1}$ $=\frac{\lambda}{n}I_{d}+\mathbb{E}_{z,\xi_{f}}\left[\frac{xx^{\top}}{d}\,\frac{d}{d\gamma}l^{\prime}\bigl{(}y,f^{\ast}_{\gamma}(x)+\gamma\bigr{)}\right],$ (68)
with
$f^{\ast}_{\gamma}=\arg\min_{f}\frac{(f-m_{f}(x)-\xi_{f})^{2}}{2\chi_{f}(x)}+l(y,f+\gamma).$ (69)
In practice, even when the data-generating distribution $p_{\rm{data}}$ is unknown, the expectations over $z$ that appear in the above expressions can be approximated by empirical averages over the observed data. This allows the generalization gap to be estimated for arbitrary distributions.
VI.2 Linear regression
We now specialize to the case of squared loss, $l(y,f)=\tfrac{1}{2}(y-f)^{2}$, with the error metric $\epsilon(y,f)=(y-f)^{2}$. In this case, the stationarity conditions take the form
$0$ $=\lambda m_{w}+n\,\mathbb{E}_{z}\left[\frac{1}{1+\chi_{f}(x)}\left(-\frac{x}{\sqrt{d}}\,(y-m_{f}(x))\right)\right],$ (70)
$\chi_{w}^{-1}q_{w}\chi_{w}^{-1}$ $=\mathbb{E}_{z}\left[\frac{xx^{\top}}{d}\,\frac{1}{(1+\chi_{f}(x))^{2}}\left((y-m_{f}(x))^{2}+q_{f}(x)\right)\right],$ (71)
$\chi_{w}^{-1}$ $=\frac{\lambda}{n}I_{d}+\mathbb{E}_{z}\left[\frac{xx^{\top}}{d}\,\frac{1}{1+\chi_{f}(x)}\right].$ (72)
The corresponding error metrics are given by
$\bar{\epsilon}_{\rm tr}$ $=\mathbb{E}_{z}\left[\frac{1}{(1+\chi_{f}(x))^{2}}\,(y-m_{f}(x)-\xi_{f})^{2}\right],$ (73)
$\bar{\epsilon}_{\rm pred}$ $=\mathbb{E}_{z}\left[(y-m_{f}(x)-\xi_{f})^{2}\right],$ (74)
$\bar{\delta\epsilon}$ $=\mathbb{E}_{z}\left[\Bigl{(}1-\frac{1}{(1+\chi_{f}(x))^{2}}\Bigr{)}(y-m_{f}(x)-\xi_{f})^{2}\right].$ (75)
In the case of linear regression with squared loss, the equation determining $m_{w}$ is independent of $q_{w}$, but it is affected by the susceptibility $\chi_{f}(x)>0$. Specifically, the effective gradient information is suppressed by the factor $1/(1+\chi_{f}(x))<1$.
VI.2.1 Remark on double descent
As discussed in Subsection V.1, understanding how fluctuations affect the extraction of information is crucial for analyzing the estimator. A particularly instructive case is the weak regularization limit $\lambda\to+0$. In this regime, the estimator reduces to the $\ell_{2}$ norm interpolator, which interpolates the training data and selects the solution of minimal $\ell_{2}$ norm in the underdetermined case. Thanks to the simplicity of the stationarity conditions in linear regression, this setting allows us to see more transparently how fluctuations affect the extraction of signal.
To this end, we assume that inputs are normalized as $\|x\|_{2}^{2}=d$ and that their components are i.i.d. It is then natural to adopt a simplified ansatz $q_{w}=\bar{q}_{w}I_{d},\;\chi_{w}=\bar{\chi}_{w}I_{d}$, so that $\chi_{f}(x)=\bar{\chi}_{w}/n$ becomes independent of $x$ (extension to more general correlated features is straightforward). Writing $\alpha=n/d$, we obtain
$\frac{1}{1+\chi_{f}(x)}=\begin{cases}1-\frac{1}{\alpha}+{\mathcal{O}}(\lambda),&d<n,\\ \frac{\lambda}{1-\alpha}+{\mathcal{O}}(\lambda^{2}),&n<d.\end{cases}$ (76)
Then, the stationarity equation for $m_{w}$ becomes
$0$ $=\mathbb{E}_{z\sim p_{\rm{data}}}\left[-\frac{x}{\sqrt{d}}(y-m_{f}(x))\right]+{\mathcal{O}}(\lambda),$ $d<n,$ (77)
$0$ $=m_{w}+\frac{n}{1-\alpha}\mathbb{E}_{z\sim p_{\rm{data}}}\left[-\frac{x}{\sqrt{d}}(y-m_{f}(x))\right]+{\mathcal{O}}(\lambda),$ $n<d.$ (78)
When $d<n$, we have $\chi_{f}(x)={\mathcal{O}}(1)$, so the signal term remains of order one in the weakly regularized limit, and $m_{w}$ approaches the ideal estimator. In contrast, when $n<d$, $\chi_{f}(x)={\mathcal{O}}(1/\lambda)$ diverges, causing the signal term to vanish at order ${\mathcal{O}}(\lambda)$ and thus introducing a bias even when $\lambda\to+0$. Although the precise form of the solution requires a more detailed analysis, the structure of the stationarity equations already reveals how such biases arise.
Similarly, the averaged generalization gap can be estimated as
$\bar{\delta\epsilon}=\overline{\rm RSS}\times\begin{cases}\frac{1}{\alpha}+\frac{1}{\alpha-1}+{\mathcal{O}}(\lambda),&d<n,\\ \frac{1}{1-\alpha}+{\mathcal{O}}(\lambda),&n<d,\end{cases}$ (79)
where $\overline{\rm RSS}=\mathbb{E}_{z\sim p_{\rm{data}}}[(y-m_{f}(x))^{2}]$. Thus, as long as the residual is nonzero, the generalization gap diverges at $d=n$, causing the double descent phenomenon. While this result is well known in solvable teacher–student scenarios Krogh and Hertz (1991); Hastie et al. (2022), the present variational framework reveals that the phenomenon can be understood without specifying a particular teacher model.
VI.3 Application to concrete problems
The most notable feature of the present formulation is that the parameters of the trial Hamiltonian can be determined adaptively for a given data-generating distribution. By approximating the population average $\mathbb{E}_{z\sim p_{\rm{data}}}$ with the empirical distribution of the observed data, the method can be applied not only to synthetic data with analytically tractable expectations, but also to real datasets where the true distribution is unknown.
In this subsection, we illustrate this point with several concrete settings. For each case, we approximate $\mathbb{E}_{z}$ by the empirical average over the observed data to determine $m_{w},q_{w},\chi_{w}$, and then evaluate the predicted generalization error $\bar{\epsilon}_{\rm pred}$. It is worth emphasizing that even with moderately large datasets, quantifying the fluctuations represented by $\chi_{w}$ and $q_{w}$ is not at all trivial without repeated experiments, yet the VGA provides a principled framework to estimate them.
(a) linear, $\lambda=10^{-2}$
(b) 2NN (relu), $\lambda=10^{-2}$
(c) 2NN (tanh), $\lambda=10^{-2}$
(d) YP MSD, $\lambda=10^{-2}$
(e) linear, $\lambda=10^{-4}$
(f) 2NN (relu), $\lambda=10^{-4}$
(g) 2NN (tanh), $\lambda=10^{-4}$
(h) YP MSD, $\lambda=10^{-4}$
Figure 1: Verification of variational Gaussian approximation (VGA) in linear regression. Markers represent true generalization errors evaluated by large test data. Lines are predictions of VGA.
VI.3.1 Synthetic data: teacher–student scenario
We first consider synthetic data in a teacher–student setting. Here the inputs are generated from a standard normal distribution, and the outputs $y$ are produced either by a linear model or by a two–layer neural network (2NN):
$x$ $\sim{\mathcal{N}}(0,I_{d}),$ (80)
$y$ $=\begin{cases}\langle x,w_{0}\rangle+\zeta,&\text{linear},\\ \tfrac{1}{\sqrt{K}}\sum_{k=1}^{K}v_{k}\,\sigma(\langle x,w_{0,k}\rangle)+\zeta,&\text{2NN},\end{cases}$ (81)
where $\sigma:\mathbb{R}\to\mathbb{R}$ denotes an activation function, $\zeta\sim{\mathcal{N}}(0,\Delta)$ is measurement noise, and the true model parameters $w_{0},v_{k},w_{0,k}$ are independently drawn from normal distributions. Throughout the synthetic cases, we set $d=100$, $K=50$, and $\Delta=0.01$.
Based on this generative model, we construct datasets of size $10^{4}$ and use them to approximate the averages over $z\sim p_{\rm{data}}$, evaluate the predicted generalization error, and compare with simulations. Note that this size $10^{4}$ is different from $n$ that appears in theoretical analysis, which is merely a control parameter. Since the components of $x$ are i.i.d., we restrict ourselves to the simplified structure $q_{w}=\bar{q}_{w}I_{d}$ and $\chi_{w}=\bar{\chi}_{w}I_{d}$.
Figure 1 (a)–(c), (e)–(g) show the comparison between the predictions of VGA and the experimental results. In all cases, the variational approximation provides a quantitatively accurate description.
VI.3.2 Real-world data
We next consider real-world data, using the Year Prediction MSD (YPMSD) dataset Bertin-Mahieux (2011). This dataset consists of the task of predicting the release year of a song from 90-dimensional audio features, with a total of 515,345515{,}345 samples.
In our setting, we assume that only 20% of the data are available, on which the VGA is performed, while the remaining 80% are hidden and used as a test set to estimate the true generalization error. All features are standardized to have zero mean and unit variance.
For the variational approximation, we consider two settings for $\chi_{w}$ and $q_{w}$: one restricted to diagonal matrices and the other allowing full matrices. The results are shown in Figure 1(d) and (h). In both cases the variational predictions are consistent with the empirical generalization errors when the full matrices are used, while the diagonal approximation exhibits deviations for small sample sizes. This indicates that correlations among features play a crucial role in this dataset.
These results demonstrate that the validity of the variational approximation extends beyond controlled synthetic settings to real-world datasets
VI.4 Single-basis ansatz
Until now, we have considered a general vector $m_{w}$ as the direction in which the parameter symmetry is broken. However, in practice one may have prior knowledge about this direction. Within the present variational framework, such prior knowledge can naturally be incorporated by restricting the structure of the trial Hamiltonian.
As the simplest example, which is particularly relevant in teacher–student scenarios or in the classification of two-component Gaussian mixtures, we consider the case where the symmetry-breaking direction is known, but its amplitude is not. Concretely, we assume
$m_{w}\propto\bar{m}_{w}w_{0},\qquad\bar{m}_{w}\in\mathbb{R},$ (82)
with a fixed reference vector $w_{0}$. The stationarity condition for $\bar{m}_{w}$ then takes the form
$\bar{m}_{w}=\frac{\frac{n}{d}\,\mathbb{E}_{z\sim p_{\rm{data}}}\left[\frac{\langle x,w_{0}\rangle\,(y-m_{f}(x))}{1+\chi_{f}(x)}\right]}{\lambda+\frac{n}{d}\,\mathbb{E}_{z\sim p_{\rm{data}}}\left[\frac{\langle x,w_{0}\rangle^{2}}{1+\chi_{f}(x)}\right]}.$ (83)
This expression makes it clear that $\bar{m}_{w}$ is determined by the (properly normalized) correlation between the residual $y-m_{f}(x)$ and the projection of the input $x$ onto the direction $w_{0}$.
In particular, consider the teacher–student model
$y=\langle x,w_{0}\rangle+\zeta,\qquad\zeta\sim\mathcal{N}(0,\Delta),$
and suppose that in the limit $d\to\infty$ the susceptibility $\chi_{f}(x)$ concentrates to a scalar $\bar{\chi}_{w}$, independent of $x$. In this case,
$\bar{m}_{w}=\frac{\frac{n/d}{1+\bar{\chi}_{w}}}{\lambda+\frac{n/d}{1+\bar{\chi}_{w}}}.$ (84)
This reproduces the well-known expression in solvable teacher–student scenario.
Although solvable settings yield compact closed forms, the variational approximation has the advantage of keeping the explicit dependence on data in the stationarity condition. This provides a clearer view of which aspects of the data determine the parameter values.
VII Summary and Conclusion
In this work, we have developed a replica analysis of inference and learning in parametric models. Unlike conventional approaches that focus on solvable teacher–student scenarios in the thermodynamic limit, we considered inference under general data distributions and finite system sizes.
To this end, we introduced a grand canonical formalism that replaces dataset averages by virtual sampling from an infinitely large data reservoir, and applied a variational approximation to the resulting replicated system. Within this framework, the stationarity conditions for the variational parameters can be derived without performing an analytic average over the data, allowing them to be determined adaptively for each given data distribution. A key technical point of our approach is that it changes the standard assumption in replica analyses: instead of requiring that the average over data can be carried out analytically, our method only requires that the trial Hamiltonian is sufficiently simple. This perspective revealed connections to well-known information criteria in statistics and machine learning. As a concrete application, we analyzed linear regression and demonstrated that the method can yield learning curves even for problems involving real-world datasets, where the data-generating distribution is not known explicitly. Although the basic idea of GC formalism was already given by Malzahn and Opper Malzahn and Opper (2002, 2000, 2001, 2005), we believe the present work provides a clearer view for analysis in parameter space.
Natural extensions of this work include analyses of more complex models, such as multilayer neural networks, for which exact solutions have not been obtained yet except for limited scenarios. Although the analysis of neural networks can become highly involved even with simple trial Hamiltonians, promising directions include structured scenarios such as sparse teacher–student models Yoshino (2020, 2023). Also, in recent years, variational approximation methods with neural networks have advanced significantly. In this work we focused on the classical variational Gaussian approximation for the sake of analytical tractability, but the development of numerical variational methods would be also an important direction for future research.
Acknowledgements.
This work was supported by JSPS KAKENHI Grant Numbers 22H05117 and 23K16960, and JST ACT-X Grant Number JPMJAX24CG.
Appendix A Grand canonical
[…]
Including this fluctuation, the replicated system may be written as
$\frac{\sum_{k=0}^{\infty}\frac{e^{-n}n^{k}}{k!}\mathbb{E}_{z\sim p_{\rm{data}}}\left[e^{-\beta\sum_{a=1}^{r}l(\theta^{a};z)}\right]^{k}e^{-\frac{\beta\lambda}{2}\sum_{a=1}^{r}\|w^{a}\|_{2}^{2}}}{\sum_{k=0}^{\infty}\frac{e^{-n}n^{k}}{k!}\,Z_{r}^{\beta,k}}$
$\propto\exp\!\left(n\,\mathbb{E}_{z\sim p_{\rm{data}}}\left[e^{-\beta\sum_{a=1}^{r}l(\theta^{a};z)}\right]-\frac{\beta\lambda}{2}\sum_{a=1}^{r}\|w^{a}\|_{2}^{2}\right).$ (85)
This expression coincides with the GC formalism  (13). From the normalization it is evident that this formulation is formally similar to a grand canonical ensemble of ideal gas in statistical mechanics, with the sample size playing the role of a fluctuating particle number and $\beta^{-1}\log n$ acting as the chemical potential. In this sense, $\tilde{D}$ may be regarded as a reservoir. This analogy provides an intuitive physical interpretation of the GC formalism. It would also be useful when considering more elaborate replica systems, such as those used in the replica analysis of bootstrap methods Takahashi (2023, 2024, 2025).
Appendix B Properties of the stationary point in a toy model
(a) $(-\partial_{\chi}{\mathcal{F}},-\partial_{q}{\mathcal{F}})$
(b) $(-\partial_{\chi}{\mathcal{F}},+\partial_{q}{\mathcal{F}})$
Figure 2: Gradient fields of the gradients of the free energy ${\mathcal{F}}\equiv\lim_{\beta\to\infty,r\to 0}{\mathcal{F}}_{r,{\rm GC}}^{\beta,n}$. (a): the steepest descent direction of $\mathcal{F}$}. (b): the case where the gradient with respect to $q$ is inverted.
In this section, by considering a simple estimation problem, we demonstrate that the stationary conditions for the trial Hamiltonian do not, in general, correspond either to maximization or minimization of the leading term of the variational free energy, $\lim_{r\to 0}r^{-1}{\mathcal{F}}_{r,{\rm GC}}^{\beta,n}$.
Let us assume that the data consist of a one-dimensional random variable $z\in\mathbb{R}$, with the loss function $l(\theta;z)=(z-\theta)^{2}/2$, parameter $\theta\in\mathbb{R}$, and $\lambda=0$. For simplicity, we also assume $\mathbb{E}_{z\sim p_{\rm{data}}}[z]=0$ and $\mathbb{E}_{z\sim p_{\rm{data}}}[z^{2}]=\Delta_{z}$. In the limit $\beta\to\infty$, the estimator $\hat{\theta}(D_{n})$ reduces to the sample mean $n^{-1}\sum_{i=1}^{n}z_{i}$. Of course, for $n\gg 1$, the central limit theorem guarantees that this converges to a Gaussian distribution with mean zero and variance $\Delta_{z}/n$ Wasserman (2013). Nevertheless, let us ignore this fact and attempt to analyze the fluctuations using the replica method.
In this case, decomposing $\hat{\theta}(D_{n})=\eta+\xi$ with $\eta\sim{\mathcal{N}}(0,\chi/(\beta n))$ representing thermal fluctuations and $\xi\sim{\mathcal{N}}(0,q/n)$ representing quenched randomness, we obtain
${\mathcal{F}}\equiv\lim_{\begin{subarray}{c}r\to 0\\ \beta\to\infty\end{subarray}}\frac{1}{\beta r}{\mathcal{F}}_{r,{\rm GC}}^{\beta,n}=\frac{q}{\chi}-\frac{n}{1+\chi/n}(\Delta_{z}+\frac{q}{n}),$ (86)
$\frac{\chi^{\ast}}{n}=\frac{1}{n-1},\quad\frac{q^{\ast}}{n}=\frac{\Delta_{z}}{n-1}.$ (87)
Thus, $\hat{\theta}(D_{n})\sim{\mathcal{N}}(0,\sqrt{\Delta_{z}/(n-1)})$, showing that the expected behavior is indeed recovered. On the other hand, this stationary point is neither a minimum nor a maximum of $\mathcal{F}$}. Figure 2 shows the gradient fields $(-\partial_{\chi}{\mathcal{F}},-\partial_{q}{\mathcal{F}})$ and $(-\partial_{\chi}{\mathcal{F}},+\partial_{q}{\mathcal{F}})$. From the figure, we see that the stationary point corresponds to a minimum in the $\chi$ direction but a maximum in the $q$ direction. Therefore, unlike in standard variational approximations, one cannot determine the trial Hamiltonian parameters simply by maximization or minimization. Unfortunately, it remains an open question which variables should be maximized and which minimized in general.